OpenStudy (ryuga):

Which expression is equivalent to this one? –7(5 + 10) A. (–7) • 5 + (–7) • 10 B. (–7) • 5 + 10 C. (–7) + 5 + 10 D. (–7) • 5 • 10

OpenStudy (ryuga):

@skullpatrol

OpenStudy (ryuga):

a?

jabez177 (jabez177):

A we know can already be ruled out. It makes no sense. Are you aware of the distributive property?

OpenStudy (ryuga):

no

jabez177 (jabez177):

Wait... You might be right. Let me re-read this. Sorry.

OpenStudy (ryuga):

who is right?

jabez177 (jabez177):

I forgot these, sorry. I'd rather let someone explain you correctly instead of me helping you get the wrong answer.

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

$$\Huge –7(5+10) = (–7)•5 + (–7)•10$$